OpenStudy (anonymous):

Match the functions to their ranges. y= 3 sin(x-pi) y=1-sin(x) y=3+4cos(x-pi) y=2+cot(x) (-infinity, infinity) [-1, 7] [-3, 3] [0, 2]

OpenStudy (xapproachesinfinity):

let's take the first one \[y=3\sin(x-\pi)\] we know that sin is always btw -1 and 1 so \[-1\le\sin(x-\pi)\leq1\] multiply by 3 \[-3\le y\le 3\] \[\therefore ~~Range ~is~~[-3,3]\]

OpenStudy (xapproachesinfinity):

repeat the same idea with others

OpenStudy (anonymous):

@xapproachesinfinity i have no idea what you did there...

OpenStudy (xapproachesinfinity):

all i did is i used our knowledge that sin and cos always trapped btw -1 an 1 ever seen the graph of sin a cos pay attention to them

OpenStudy (xapproachesinfinity):

then i multiplied by 3 so i get my functions range

OpenStudy (anonymous):

do i always multiple by 3. ?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

no depends on the function you have that one has 3 to it but other will be different

OpenStudy (anonymous):

okay, then i don't know what im doing. so im going to guess. lol